OpenStudy (anonymous):

HELPPP justin buys a motorcycle for $10,000. The dealer is charging him an annual interest rate of 9.25% and is using the add-on method to compute monthly payments. If he pays off the loan in 72 months, what are his monthly payments? If he makes a down payment of $2,900, how much will his monthly payments be? Do not round until the final answer. Then, round to the nearest cent. A$432.78; $153.34 B $215.97; $279.14 C $215.97; $153.34 D $215.97; $307.27

OpenStudy (anonymous):

The interest formula would apply here, p=c(1+r)^t

OpenStudy (anonymous):

p=10,000(1.00925)^6

OpenStudy (anonymous):

9%=0.09 for reference

OpenStudy (anonymous):

And excuse me, this would be exponential decay. So let me start from scratch.

OpenStudy (anonymous):

So p=c(1-r)^t

OpenStudy (anonymous):

p=10,000(1-0.0925)^6 72 months is equal to 6 years. Annual=Yearly

OpenStudy (anonymous):

p=10,000(0.9075)^6

OpenStudy (anonymous):

so 5585.72?

OpenStudy (anonymous):

oh well no cuz thats not in the answers above , so im not surewhat im doing wrong

OpenStudy (anonymous):

I'll explain it.

OpenStudy (anonymous):

First, we need to find out how much the interest is for each year. He is also paying monthly payments, not annual.

OpenStudy (anonymous):

oh ok i think i got the first part so its 215.97

OpenStudy (anonymous):

and then for the second part i minus 2900 from 10000

OpenStudy (anonymous):

So, the best thing to do first is to find the interest added each year. After we do that, we will be able to determine the how much he had to pay off each month in order to be able to pay off the 10K in 72 months.

OpenStudy (anonymous):

p=7100(0.9075)^6

OpenStudy (anonymous):

Let's find 0.9075^6 first

OpenStudy (anonymous):

oh ok i think i understand now

OpenStudy (anonymous):

0.558572823884 is the answer, but rounded, it is 0.56

OpenStudy (anonymous):

7100(0.56)=P, interest added

OpenStudy (anonymous):

P=3976, which is the interest added if he were to add a down payment first.

OpenStudy (anonymous):

ok thank you so much for your help

OpenStudy (anonymous):

i have another question im not sure if you can help

OpenStudy (anonymous):

And to what you said, 5585 is the interest for the entire 10K

OpenStudy (anonymous):

So it would be 15585 for the first one.

OpenStudy (anonymous):

A credit card calculates interest using the average daily balance method. The card charges 20.5% annual interest rate on the average daily balance. The following transactions occurred during the March 1 – March 31 billing period. Find the average daily balance for the billing period. Round to the nearest cent.

OpenStudy (anonymous):

Okay this one,

OpenStudy (anonymous):

You have to add up all of the transactions. The Credit transaction is a Negative, since he is borrowing money.

OpenStudy (anonymous):

His starting balance is 2653.49

OpenStudy (anonymous):

so thats the average daily balance ?

OpenStudy (anonymous):

After you add up all of the numbers, you will get about 796.71

OpenStudy (anonymous):

June has 30 days in it, so divide this by 30.

OpenStudy (anonymous):

divide 796.71 by 30 ?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

26.557

OpenStudy (anonymous):

Correct, round the 57 to the nearest hundreth, since we don't get anything lower than pennies.

OpenStudy (anonymous):

.557 rounded to the nearest hundreth is .56, therefore the daily balance is 26.56

OpenStudy (anonymous):

ok thank you i have such a hard time with these and i have like 7 more

OpenStudy (anonymous):

thansk you helped alot